OpenStudy (anonymous):

parallel to y=3x+5 and passes through (2,-1)

OpenStudy (campbell_st):

what is the slope of the given line...?

OpenStudy (anonymous):

3

OpenStudy (jonnyvonny):

So for something to be parallel, it has to have the same slope, the number tagged along with the "x". So, the slope of this function is "3". The only difference, then, since its the same line, is where it starts, which is the y-intercept. -1=3(2)+b -1=6+b -7=b Now, you have both the slope of the new graph, which is 3, and the y-intercept, which is -7. Now make the equation.

OpenStudy (campbell_st):

that's correct... the line is \[y = 3x + b\] to find b, substitute x = 2 and y = -1 and solve for b, the y-intercept

OpenStudy (anonymous):

is it y=3x-7?

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

thank you